If you wish to accumulate $200,000 in the child's college fund, and can invest at a 7.5% annual rate, how much must you invest at the end of each year if the first deposit is made at the end of the first year (round to nearest $1)?
Answer #1

Investment at the end of each year=F*i/[(1+i)^n-1]

F is the future value
i is the annual rate
n is the number of years

Now, substituting the values in the equation, we get;
=200000*7.5%/[(1+7.5%)^18-1]
=15000/2.675804089
=5605.791568 or $5606 (Rounded to nearest dollar)
So, the investment at the end of each year is $5606